How would you solve y=-2x+3 and x+y=4? Using the Substitution method.
Answer by TimothyLamb(4379)

y=-2x+3
x+y=4
---
put the system of linear equations into standard form
---
2x + y = 3
x + y = 4
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -1
y = 5
---
now solve by substitution:
---
x + y = 4
x - 2x + 3 = 4
-x = 1
x = -1
y = -2(-1) + 3
y = 2 + 3
y = 5
---
solution by substitution:
x = -1
y = 5
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---